### 31 Jan 2024 v1

# SeFi-IDE: Semantic-Fidelity Identity Embedding for Personalized Diffusion-Based Generation

Abstract
00footnotetext: $*{}^{*}$ indicates equal contribution.00footnotetext: indicates corresponding author.
Advanced diffusion-based Text-to-Image (T2I) models, such as the Stable Diffusion Model, have made significant progress in generating diverse and high-quality images using text prompts alone. However, T2I models are unable to accurately map identities (IDs) when non-famous users require personalized image generation. The main problem is that existing T2I models do not learn the ID-image alignments of new users. The previous methods either failed to accurately fit the face region or lost the interactive generative ability with other existing concepts in T2I models (i.e., unable to generate other concepts described in given prompts such as scenes, actions, and facial attributes). In this paper, we focus on accurate and semantic-fidelity ID embedding into the Stable Diffusion Model for personalized generation. We address this challenge from two perspectives: face-wise region fitting, and semantic-fidelity token optimization. Specifically, we first visualize the attention overfit problem, and propose a face-wise attention loss to fit the face region instead of the whole target image. This key trick significantly enhances the ID accuracy and interactive generative ability with other existing concepts. Then, we optimize one ID representation as multiple per-stage tokens where each token contains two disentangled features. This expansion of the textual conditioning space enhances semantic-fidelity control. Extensive experiments validate that our results exhibit superior ID accuracy and manipulation ability compared to previous methods. More results: https://com-vis.github.io/SeFi-IDE/.
Figure 1: The face identity (ID) embedding of previous personalized generation methods has two problems:(1) Attention Overfit : The attention of Textural Inversion [8] and ProSpect [48] towards to the “$*V*$” takes over the whole image, disrupting the interaction between newly embedded concepts and other existing concepts. Consequently, their attention towards to the “cup” is wrong, which results in the failure of the given prompt. (2) Limited Semantic-Fidelity: Despite alleviating overfit, Celeb Basis [47] introduces excessive face prior, limiting the semantic-fidelity of the learned ID embedding. Specifically, this means that the “cup” attention map continues to concentrate on the face region. We propose Face-Wise Attention Loss (Sec. 3.1) and Semantic-Fidelity Token Optimization (Sec. 3.2) to address problem (1) and (2) respectively.
1 Introduction
Recently, Text-2-Image (T2I) models, such as the Stable Diffusion Model [29], have demonstrated an impressive ability to generate diverse, high-quality, and semantic-fidelity images using text prompts alone, thanks to image-aligned language encoders [25] and diffusion-based generative models [5, 21]. However, the challenge of personalized generation still remains, because the accurate person-specific face manifold can not be represented by text tokens, especially for the non-famous users whose data are not included in the training dataset. In this paper, we focus on accurate and semantic-fidelity identity embedding for personalized diffusion-based generation using one face image.
The previous methods have two problems require to be addressed: (1) Their fine-tuning strategies [8, 19, 30] tend to fit the whole target image rather than the face region. This resulted in the low ID accuracy and the difficulty of generating other concepts in the given prompt, such as ID-unrelated scenes (e.g., Eiffel Tower), ID-related facial attributes (e.g., expressions and age), and actions (e.g., playing the guitar). Particularly for actions, it is more challenging to generate prompt-fidelity human motions and human-object interactions, which can be shown in Fig. 1. (2) Their ID embedding methods lack the semantic-fidelity representations for facial attributes, which results in that human faces are treated as objects without non-rigid and diverse deformations. Although the method such as Celeb Basis [47] can achieve an accurate ID mapping, it is unable to manipulate the facial attributes of the target image, such as expressions.
To address these problems, we propose our identity embedding method from two perspectives: (1) Face-Wise Region Fit: We first visualize the attention overfit problem of the previous methods from the attention feature activation maps, and then propose a face-wise attention loss to fit the face region instead of the whole target image. This key trick can improve the ID accuracy and interactive generative ability with the existing concepts in the original Stable Diffusion Model. (2) Semantic-Fidelity Token Optimization: We optimize one ID representation as several per-stage tokens and each token consists of two disentangled features. This approach expands the textual conditioning space and allows for semantic-fidelity control ability. Our extensive experiments confirm that our method achieves higher accuracy in ID embedding and produces a wider range of scenes, facial attributes, and actions compared to previous methods.
To summarize, the contributions of our approach are:
  • We visualize attention overfit problem of the previous methods, and propose a face-wise attention loss for improving the ID embedding accuracy and interactive generative ability with the existing concepts in the original Stable Diffusion Model.
  • For semantic-fidelity generation, we optimize one ID representation as several per-stage tokens with disentangled features, which expands the textual conditioning space of the diffusion model with control ability for various scenes, facial attributes, and actions.
  • Extensive experiments validate our advantages in ID accuracy and manipulation ability over previous methods.
  • Moreover, our embedding method has the potential to be applied to other categories, including dogs, due to its independence from any prior facial knowledge.
Figure 2: The overview of our framework. We first propose a novel Face-Wise Attention Loss (Sec. 3.1) to alleviate the attention overfit problem and make the ID embedding focus on the face region to improve the ID accuracy and interactive generative ability. Then, we optimize the target ID embedding as five per-stage tokens with disentangled features to expend textural conditioning space with semantic-fidelity control ability (Sec. 3.2).
2 Related Work
2.1 Text-Based Image Synthesis and Manipulation
Synthesis. Previous models such as GAN [16, 10, 45, 44, 11, 43], VAE [18, 37, 46, 42], Autoregressive [27, 7], Flow [26, 6] were adopted to model the dataset distribution. Recent advanced diffusion models [5, 21] have shown excellent diversity and fidelity in text-to-image synthesis [22, 14, 27, 29, 31, 2]. Conditioned on the text embedding of the text encoder [25], these diffusion-based models are optimized by a simple denoising loss and can generate a new image by sampling Gaussian noise and a text prompt.
Manipulation. However, these diffusion-based generative models have difficulties in generating novel concepts [3] or controlling the identity of generated objects [8]. Existing methods either directly blended the latent code of objects [1] to the generated background, or failed to understand the scenes correctly [41], which results in the obvious artifacts. To further solve this problem, some works [12, 20, 23, 24] adopted attention-based methods to manipulate target objects, but cannot balance the trade-off between content diversity and identity accuracy.
2.2 Personalized Generation of T2I Models
The existing diffusion-based T2I models are limited by generating ID-accurate images for the specific unseen concepts. Thus, using images from the novel concepts for fine-tuning enables us to obtain a personalized model, which can insert novel concepts into the original model and synthesize ID-specific new scenes, appearances, and actions. Inspired by the GAN Inversion [39], recent diffusion-based personalized generation works can be divided into two categories: (1) Fine-Tuning T2I model: DreamBooth [30] fine-tunes all weight of the T2I model on a set of images with the same ID and marks it as the specific token. (2) Token Optimization: The works such as Textural Inversion [8], ProSpect [48], and Celeb Basis [47] optimize the text embedding of special tokens to map the specific ID into the T2I model, where the T2I model is fixed in the optimization process. Recently, these two strategies are combined for manipulating multi-concept interactions [19] or saving fine-tuning time and parameter amount [33, 35].
ID Embedding for Faces. Previous methods [32, 9] try to train a inversion encoder for face embedding, but face ID-oriented mapping is difficult to be obtained from a naively optimized encoder. Moreover, fine-tuning the T2I model on large-scale images often causes concept forgetting. Towards this, Celeb Basis [47] adopt a pre-trained face recognition model and a face ID basis to obtain an ID representation for one single face image and Face0 [36] learned to project the embeddings of recognition models to context space of Stable Diffusion. However, introducing too strong face prior makes it difficult to manipulate diverse facial attributes and fails to generalize to other concept embedding.
3 Method
Embedding one specific ID into the Stable Diffusion Model for personalized generation using only one single face image has two technical requirements: accuracy and semantic-fidelity. As shown in Fig. 2, we propose our ID embedding pipeline from two key perspectives: (1) Face-Wise Attention Loss: Towards the improvements of ID accuracy and interactive generative ability with the existing concepts in the original model, we propose a face-wise attention loss in Sec. 3.1. (2) Semantic-Fidelity Token Optimization: For diverse manipulation, we optimize one ID representation as several per-stage tokens, and each token consists of two disentangled features, which can be seen in Sec. 3.2.
3.1 Face-Wise Attention Loss
We first analyze and visualize the attention fit problem of previous methods. Then, we present an accessible prior from the Stable Diffusion Model to improve the interactive generative ability.
Image Fit vs. Face-Region Fit. Previous methods rely on learning from multiple images of a target object to grasp relevant features, such as Textual Inversion [8] and DreamBooth [30]. However, when only a single image is available, they are prone to fitting to the whole target image and the learned embedding tokens tend to influence regions beyond the face region during the cross-attention stages. As a result, they lack the good interactive generative ability with the existing concepts in the original model [35]. In other words, during the inference, the generated results from the personalized model may not be consistent to the text prompts. For example, as shown in Fig. 1, the given prompt is “a V* is enjoying a cup of latte”, but the methods with attention overfit problem fail to generate the “cup” content. The same problem can also be seen in Fig. 3, which given some facial attribute such as “old” in the prompt, this diffusion-based generation process just fails. Our ID embedding optimization can focus on the ID-related face regions and neglect the unrelated background features, which can improve the ID accuracy and interactive generative ability.
Make Best of Stable Diffusion Model. Multiple target images are necessary for previous methods to acquire concept-related embedding tokens. These images enable users to utilize text prompts to manipulate different object poses, appearances, and contextual variations. One target image fails to achieve such generalization. However, Stable Diffusion Model has learnt a strong general concept prior for various sub-categories. For example, different human identities belong to the general concept “person”, and different dog categories such as Corgi and Chihuahua belong to the general concept “dog”. Therefore, it is reasonable to adopt this prior knowledge to achieve one-shot learning. To meet our more higher requirements, we aim to have flexibility to manipulate the ID-specific regions of final images. In other words, when we want to generate images corresponding to “a photo of V is playing guitar”, only handling portrait or face image generation is not enough for this prompt. Therefore, we adopt “the face of person” as our general concept prior for ID embedding, because when provided with the prompt “a photo of the face of person”, Stable Diffusion Model can generate an image of a person with a randomly assigned identity and constrain the region where the generated person appears in the final image.
We propose to use a reference prompt $P_{r}$ that remains consistent with the general concept of different IDs, which replaces the placeholder word (“V”) with “person” in “a photo of the face of V”. Then, we use this attention map derived from $P_{r}$ as a constraint to restrict the attention corresponding to the placeholder word (“V*”) in the target prompt “a photo of the face of V*”. This approach allows the ID embedding to focus on the face region associated with the target ID while maintaining the coherence of the general concept. Specifically, we first utilize CLIP encoder [31] to encode the reference prompt and target prompt as $\bm{P_{r}}$ and $\bm{P_{t}}$. The $\bm{P_{t}}$ is fed to the self-attention module to obtain five per-stage tokens $\bm{P}={\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$. Then, we adopt text encoder [31] to obtain their corresponding key ($\bm{K}$) and value ($\bm{V}$) features ${\bm{K_{r}},\bm{V_{r}}}$ and ${\bm{K_{t}},\bm{V_{t}}}$. Then, each of the K features are send to the cross-attention module to calculate the attention map $\bm{A_{r}}$ and $\bm{A_{t}}$ with latent image features $\bm{Q}$ respectively. The $\bm{A_{t}}$ is constrained by the $\bm{A_{r}}$ within the corresponding concept representation as follows:
$\mathcal{L}_{attention}=||\bm{A_{r}}-\bm{A_{t}}||_{2}^{2}$ (1)
3.2 Semantic-Fidelity Token Optimization
We first present the disadvantages of previous methods. And then we introduce our optimization strategy.
Lack of Semantic-Fidelity Control. This problem can be found from two perspectives: (1) Stable Diffusion Model: We observe that even though the face data of celebrities has been included in the training dataset of Stable Diffusion Model, it still fails to achieve perfect semantic-fidelity control for these IDs. For example, “a photo of an old Obama” can not generate corresponding images. (2) Previous Personalized Methods: These works [48, 47] based on Textural Inversion [8] mainly emphasize how to preserve the characteristics of the person and achieve global control over the generated images through text modifications. Although these methods are able to manipulate scenes or styles, they struggle to control fine-grained facial attributes of learned IDs, such as age and expressions.
Disentanglement of $\bm{K}$ and $\bm{V}$ Features. The text prompt will be encoded into text features ($\bm{K}$, $\bm{V}$) by the clip text encoder of Stable Diffusion Model, and then fed into cross-attention layers of U-Nets for conditioning the generated images. Previous methods [19, 35] differentiated the $\bm{K}$ and $\bm{V}$ matrices as position information and object texture features is not appropriate for control facial attributes. As shown in Fig. 5, we found that $\bm{K}$ and $\bm{V}$ features have different semantic meanings. The $\bm{V}$ feature is more ID-related, while the $\bm{K}$ feature is more related to environment factors such as lighting. The disentangled optimization of $\bm{K}$ and $\bm{V}$ features can further improve the ID accuracy and interactive generative ability.
Specifically, the input prompt is fed into the CLIP text encoder, which generates the textual token embedding $\bm{P_{t}}$. The $\bm{P_{t}}$ is fed to a self-attention [38] module, which consists of two self-attention layers with one feed-forward layers. Finally, we evenly divide the 1000 diffusion steps into five stages and optimize one ID embedding as five per-stage tokens $\bm{P}={\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$, while each token consists of a $\bm{V}$ feature and a $\bm{K}$ feature as follows:
$\mathcal{L}_{K,V}=\mathbb{E}_{\bm{z},\bm{t},\bm{P}}[||\epsilon-\epsilon_{\theta}(\bm{z_{t}},\bm{t},(\bm{P_{i}^{K}},\bm{P_{i}^{V}}))||_{2}^{2}],$ (2)
where $\bm{z}$ is the latent code of target image . The $\epsilon$ is the unscaled noise sample, and the $\epsilon_{\theta}$ is the U-Net modules in diffusion Model. The $\bm{t}$ is the optimization step of the diffusion process, and the $\bm{z_{t}}$ is U-Net output of different steps.
Our goal is seamlessly embedding one specific ID into the space of Stable Diffusion Model, which have to achieve accurate ID mapping and fully use the prior from the Stable Diffusion Model to manipulate scenes, actions and facial attributes. Thus, the total optimization objective can be formulated as follows:
$\vspace{-0.4cm}\mathcal{L}=\mathcal{L}_{K,V}+\lambda\mathcal{L}_{attention}.$ (3)
Figure 3: Qualitative comparisons with different SOTA methods using different prompts.
4 Experiments
4.1 Experimental Settings
Implementation Details. (1) Pre-Trained Model: In all experiments, we utilize Stable Diffusion 1.4 [29] with default hyper parameters as the pre-trained diffusion-based T2I model. We adopt a frozen CLIP model [31] in the Stable Diffusion Model as the text encoder network. The texts are tokenized into start-token, end-token, and 75 non-text padding tokens. (2) Test Data: The test face images are selected from the CelebA-HQ dataset [15] and then resized to 512. (3) Training Details: For our method, the time for fine-tuning every ID using only one face image is $\sim 5$ minutes ($\sim 1000$ epochs) on one NVIDIA TITAN RTX GPU. We adopt Adam optimizer and set its learning rate as 0.005. The $\lambda$ is set as 0.003.
Baselines. Our task setting is using only a face image to embed ID into the Stable Diffusion Model. Thus, for fair comparisons, we only use a single image for all personalized generation methods but using enough optimization time for different methods. We select five state-of-the-art works as baseline methods for comparisons from two perspectives: (1) Model Fine-Tuning: DreamBooth [30] (learns a unique identifier and fine-tunes the diffusion model to learn from a set of images). Custom Diffusion [19] (retrieves images with similar captions of the target concept and optimizes the cross-attention module with a modifier token), (2) Token Optimization: Textual Inversion [8] (learns a pseudo-word for a concept within a limited number of images for optimization), ProSpect [48] (expands the textual conditioning space with several per-stage textual token embeddings), and Celeb Basis [47] (builds a well-defined face basis module to constrict the face manifold).
Figure 4: The manipulation diversity of our methods, which is shown from stylization, action, and facial attribute.
[…]
Prompt ID Detect ID (Prompt) Prompt ID Quality Train Add (min)
[…]
Table 1: Quantitative evaluation between different SOTA methods and ours.
Metrics. Our quantitative evaluation is conducted from objective metrics, user study, parameter amount, and fine-tuning time. (1) Objective Metrics: We select Prompt (CLIP score [31]), ID (ID feature [4] similarity), and Detect (face detection rate [4]). Furthermore, to reflect the essence of T2I generation and fair ID accuracy evaluation, we propose a new metric for face personalized generation which is noted as ID (Prompt). Specifically, if the CLIP score of this image is lower than the threshold (set as 0.23), then the ID (Prompt) score of this image is 0. The threshold is the average CLIP score of these images which get higher scores in user study. (2) User Study: We select more than $\sim$20 volunteers and generate $\sim$200 images, to evaluate different methods from prompt-image consistency (Prompt), ID accuracy (ID), and image quality (Quality). (3) Parameter Amount: We compare the parameter amount from parameters to be learned (Train) and the total introduced parameters (Add). (4) Time: We evaluate the fine-tuning time.
Figure 5: The different effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens.
Figure 6: Different options for attention loss calculation.
4.2 Comparisons with State-of-the-Art Methods
Qualitative Evaluation. We utilize the same prompt to test five different state-of-the-art personalized generation methods and ours, the qualitative results are shown in Fig. 3. We evaluate the performance from three different levels: facial attributes (e.g., age and hairstyle), actions (i.e., human motion and interactions with other objects), and combinations of facial attributes and actions. Textural Inverion [8] and Prospect [48] tend to overfit the input image, so they fail to interaction with other concepts. Although DreamBooth [30], Custom Diffusion [19] successfully generate the interaction image of human and concept, the generated identities fail to maintain the consistency with the target images. Celeb Basis [47] successfully generate the human-object interaction actions, but they fail to manipulate the facial attributes of target identities well.
[…]
Table 2: Quantitative evaluation of ablation study.
Quantitative Evaluation. As shown in Tab. 1, our method achieves the SOTA performance in the image-prompt alignment evaluation, and achieves similar face detection rate to Textural Inversion [8]. Due to attention overfit, although Textural Inversion [8]’s ID and Detect evaluation scores are higher than the other methods, it fails to handle the interactive generation between new ID embedding and existing concepts in the original model. For fair ID evaluation, we propose a new metric, which is noted as ID (Prompt), which requires the generated images have to achieve the task of semantic-fidelity, then we calculate their ID scores. Our method achieve better ID (Prompt) scores than Textural Inversion [8] and ours is excellent in image-prompt alignment evaluation, and this improvement is from two reasons: (1) Attention Overfit Alleviation: our face-wise attention loss is able to alleviate the attention overfit problem of previous methods such as DreamBooth [30], Prospect [48], and Texutral Inversion [8]. Our method can make the ID embedding focus on the face region, instead of the whole image. (2) Attribute-Aware Tokens: Compared to Celeb Basis [47], our method does not introduce too much face prior and represent one ID as five per-stage tokens, which can balance the trade-off between ID accuracy and manipulation ability. Our expended textural conditioning space have a strong disentanglement and control ability of attributes (e.g., action-related objects and facial attributes) than Celeb Basis [47].
Figure 7: Utilizing different number of K-V pairs.
User Study. To make our results more convincing, we further conduct a user study, which can be found in Tab. 1. Our method obtains better preference than previous works among the participating users.
Figure 8: Multi-ID embedding and generation.
Efficiency Evaluation: Parameter Amount and Time. As shown in Tab. 1, we have advantages in introduced parameter amount and fine-tuning time. The Celeb Basis [47] introduces a basis module and a pre-trained face recognition model, but these are big optimization burdens and a too strong facial prior can disrupt the interaction between faces and other concepts. We utilize the prior from the T2I model itself to embed IDs, reducing the introduction of additional parameters and further enhancing the facial manipulation ability of the T2I model.
4.3 Ablation Study
Different Effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ Tokens. The semantic information of per-stage tokens is important for the interpretation of diffusion-based generation process, especially for the different effects of $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens. As shown in Fig. 5, we thoroughly explore this from two perspectives: (1) Progressively Adding: We add different ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$ tokens to the conditioning information in ten steps. We found that the initial tokens effect more the layout of generation content (e.g., face region location, and poses), while the latter tokens effect more the ID-related details. (2) Progressively Substituting: We then substitute different $\bm{P_{i}^{K}}$ and $\bm{P_{i}^{V}}$ tokens of ${\{(\bm{P_{i}^{K}},\bm{P_{i}^{V})}\}}_{1\leq i\leq 5}$. We found that $\bm{P_{i}^{V}}$ contribute to the vast majority of ID-related conditioning information, and the $\bm{P_{i}^{K}}$ contribute more to textural details, such as environment lighting.
Figure 9: Using our ID embedding method for non-face concepts.
Attention Loss. We thoroughly investigate three options for face-wise attention loss. The option #$\#1$ only regularizes on the $*V*$ token and the option #$\#2$ regularizes the prompt-length tokens. As shown in Fig. 6, option #$\#1$ effects the other concept embeddings in the T2I model, which results in non-ID concepts that cannot be generated, such as sunglasses. Although the option #$\#2$ can reduce the influence of too much ID attention, the activation region of $*V*$ still disrupts regions beyond its scope, which can be seen in the corners of the feature activation map for $*V*$. Our final adopted option is #$\#3$, which calculates the attention loss among the whole text attention maps generated by each token. This option prevents the learned token from overfitting to other regions and only focus on the face region.
The Number of K-V Pairs. As shown in Fig. 7, we explore the influence of K-V pair numbers. When using only one pair of K-V, the learned ID-related tokens fail to maintain good ID accuracy and interact with other complex concepts and attributes. However, adopting too many K-V pairs (e.g., 10 pairs) fails to bring significant improvements of diversity or quality, and this is no doubt a huge computational burden. In our method, we select 5 K-V pairs, which balances the trade-off of representing capacity and computation.
4.4 Generalization: Multi-ID and Other Objects
We further validate our methods on generation of multiple IDs. The results in Fig. 8 show that training individually can also perform good performance of human-human interactions. Furthermore, compared to Celeb Basis [47], our method does not introduce face prior from other models (e.g., a pre-trained face recognition model or basis module). As shown in Fig. 9, we adopt bear, cat, and dog for experiments, which show the generalization ability of our method.
5 Conclusion
We propose two novel problem-orient techniques to enhance the semantic-fidelity of ID embedding for personalized diffusion-based generation. We analyze the attention overfit problem and propose Face-Wise Attention Loss. This improves the ID accuracy and facilitates the effective interactions between this ID embedding and other concepts (e.g., scenes, facial attributes, and actions). Then, we optimize one ID embedding as multiple per-stage tokens, which further expands the textual conditioning space with semantic-fidelity control ability. Extensive experiments validate our better ID accuracy and manipulation ability than previous methods, and we thoroughly conduct ablation study to validate the effectiveness of our methods.